A constant force of magnitude 50N is applied ona body of mass 10kg moving initially with a speed of 10 m s-1 against the direction of motion. how long will it take the body to come to rest? 

Grade:12

1 Answers

Yadesh
5 years ago
Retardation produced in body is ==Force÷mass=50÷10=5m/s^2 
So now by eqn--
v=u+at
0=10+(--5)t
Or 10=5t
Hence time=2sec is answer